# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ Appendix I Noise amplitude in the transformed reward estimate
offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1376}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1397}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1368}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1392}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ 1e-4 3e-7 100
offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1426}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1444}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1426}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1442}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 1e-4 1e-6 100
@@
SFT $\rightarrow$ offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1398}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0000}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1409}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1398}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0000}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1408}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 1e-4 1e-7 200
SFT $\rightarrow$ offpolicy–random* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1437}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1457}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1445}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0008}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1466}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ 1e-4 1e-6 100
@@
offpolicy–best* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1458}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1469}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1460}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1470}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ 0.003 1e-6 200
offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1335}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1349}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0009}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1333}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1348}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0009}}$ 0.001 1e-6 0

### v2

@@ Appendix I Noise amplitude in the partition function estimate for the initial reward
off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1376}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1397}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1368}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1392}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ 1e-4 3e-7 100
off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1426}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1444}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1426}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1442}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 1e-4 1e-6 100
@@
SFT $\rightarrow$ off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1398}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0000}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1409}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1398}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0000}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1408}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 1e-4 1e-7 200
SFT $\rightarrow$ off-policy–random* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1437}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1457}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1445}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0008}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1466}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ 1e-4 1e-6 100
@@
off-policy–best* ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1458}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1469}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\underline{0.1460}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1470}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ 0.003 1e-6 200
off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1335}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1349}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0009}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1333}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1348}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0009}}$ 0.001 1e-6 0